1. If your effective annual discount rate is 7%, what is your effective quarterly interest rate? (express as a decimal to 3 digits, e.g. 7% should be written as 0.070)

2. If you had to amortize a $120,000 loan over a 10-year period into a payment stream that looks like a uniform annuity flow taking the time value of money into an account, then the value of the monthly payments at an APR of 9% on the loan must be how much? (round to the nearest whole dollar)

3. Epsilon Enterprises announces that it will pay a dividend of $2 next period to its shareholders. If the dividends are expected to grow at a rate of 1.2%, what is the price of a share of stock predicted by the dividend discount model of equity valuation at an opportunity cost of 6.3%? (round to the nearest whole dollar)

4.

Mr. Peanutbutter and Princess Buttercup are both evaluating the same cash flow stream. Princess Buttercup, however, has better outside investment opportunities than Mr. Peanutbutter does and she faces a higher opportunity cost of time as a result. Which of the following will be true?

a) Buttercup and Peanutbutter will have identical willingness to pay for the same cash flow stream

b) Buttercup will be willing to pay more than Peanutbutter for the same cash flow stream

c) Not enough information

d) Peanutbutter will be willing to pay more than Buttercup for the same cash flow stream

5.

If a bond's coupon rate is smaller than the yield to maturity, then

a) Not enough information

b) The face value payment must equal the initial purchase cost

c) The face value payment must exceed the initial purchase cost

d) The face value payment must fall short of the initial purchase cost

6.

Which of the following bonds is different than the others?

a) (-90, 15, 15, 15, 105)

b) (-18, 3, 3, 3, 21)

c) (-45, 7.5, 7.5, 7.5, 50)

d) (-6, 1, 1, 1, 7)

7.

When the central bank raises the interest rates, then generally

a) bond prices decrease and stock prices increase

b) bond prices and stock prices tend to decrease

c) bond prices and stock prices tend to increase

d) bond prices increase and stock prices decrease

8.

Which of the following cash flow streams represent bonds that are trading at par? (select all that apply)

a) (-50, 2, 2, 50)

b) (-100, 5, 5, 105)

c) (-10, 1, 1, 12)

d) (-20, 1, 1, 21)

9.

Which of the following must be true regarding the bond described by the cash flow stream (-100, 5, 105)? (select all that apply)

a) the bond is trading at par

b) it is equivalent to the bond (-50, 3, 53)

c) the coupon rate is 5%

d) the yield on the bond is 5%

10.

If the coupon rate on a bond is greater than the yield on the bond, then which of the following must be true? (select all that apply)

a) The face value exceeds the price of the bond

b) The bond is trading above par

c) The price of the bond exceeds the face value

d) The bond is trading under par
Answer #1

Solution :- (1)

Effective Annual Interest Rate = 7.00%

Effective Quarterly Interest Rate = ( 1 + 0.07 )1/4 - 1

= 1.01706 - 1

= 0.01706

= 1.706%

Solution :- (2)

Amount Financed = $120,000

Total Monthly Payments = 10 * 12 = 120

Interest Rate per month = 9% / 12 = 0.75%

Now Value of Monthly Installment = $120,000 / PVAF ( 0.75% , 120 )

= $120,000 / 78.9417

= $1,520.11

The Value of Monthly Payment = $1,520.11

Solution :- (3)

Price of a Stock = D1 / ( ke - g )

= $2 / ( 0.063 - 0.012 )

= $39.22

if there is any doubt please ask in comments

Thank you please rate

As per HomeworkLib policy we need to answer only one question at once even i answer three so please ask others as seperate one